OpenStudy (anonymous):

Given f(x) = 3x – 5, what is f(–4)? A. B. C. D.

OpenStudy (anonymous):

3(-4) - 5 = -12-5 = -17

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcme :)